A boy of mass 40kg runs up a flight of 30 steps each 150mm high in 6secs. Find the average power develop by the boy (take g=10m/s^2)


1
Expert's answer
2021-12-03T12:44:55-0500
"P=\dfrac{W}{t}=\dfrac{mgnh}{t},""P=\dfrac{40\ kg\times10\ \dfrac{m}{s^2}\times30\times0.15\ m}{6\ s}=300\ W."